what is the soultion to this simultaneous equation using substion or addtion method?
x-y=0
7x-3y=24

Answer by sarah_adam(201)

x-y=0 --- eq 1
7x-3y=24 --- eq 2
3x - 3y = 0 --- eq1 *3 (myultipling equation 1 with 3)
7x - 3y = 24 --- eq2
(-) (+) (-)
-------------
-4x = - 24 (subtracting eq1 and eq2)
x = -24/-4 = 6
Substitute value of x in eq1
x - y = 0
y = x
y = 6
x = 6; y = 6